OpenStudy (hockeychick23):

What is the probability that out of 125 babies born, at least 60 will be girls? Assume that boys and girls are equally probable. (I got the answer 70.54 I just want to make sure I did it correctly, thanks!)

OpenStudy (perl):

we can use the formula we used in the previous problem

OpenStudy (perl):

we want probability 60 or greater

OpenStudy (hockeychick23):

ok thanks, what i did this time was i did 125(.5)= 62.5 sqrt(31.25) then: z=60-62.5/sqrt(31.25)= -.4472135955 and then looked it up on the z score chart and thats how i got the answer

OpenStudy (perl):

right, just make sure you know what type of z table it is sometimes it uses P ( Z < a ) , sometimes it uses P ( Z > a) if it is the latter, you will need to subtract by 1

OpenStudy (hockeychick23):

ohh ok so is that where i messed up?

OpenStudy (hockeychick23):

also this is where i got the table from http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf

OpenStudy (perl):

yes that table gives you P ( Z < a )

OpenStudy (perl):

also you need to make a binomial continuity correction

OpenStudy (perl):

we are going to find P( X >= 59.5 ) , since X = 60 is a rectangle from 59.5 to 60.5

OpenStudy (hockeychick23):

ok cool so do i just replace the 60 with 59.5 in the same equation?

OpenStudy (perl):

yes, to find the z score

OpenStudy (perl):

it turns out on calculator normalcdf ( 60, 1E99, 125*.5 , sqrt( 125 * .5 * .5) ) = .6726 , which is wrong

OpenStudy (perl):

thats why we need the 'binomial correction factor '

OpenStudy (hockeychick23):

yea i replaced it at 59.5-62.5/5.59 and got 3/5.59= .5366726

OpenStudy (perl):

that should be -.5367

OpenStudy (perl):

and now P( Z > -.5367) = 1 - P ( Z < -.5367) by complement rule

OpenStudy (hockeychick23):

ok, so i shouldn't use the z-score table but ue the complement rule instead?

OpenStudy (perl):

youre doing it correct

OpenStudy (perl):

we just have to use complement rule, since the table only shows one direction

OpenStudy (perl):

lets round -.5367 to -.54 P( Z < -.54) = .2981 now we want P( Z > -.54) but that is equal to 1 - P ( Z < -.54) = 1- .2946 this is only accurate to 2 decimal places i would say

OpenStudy (hockeychick23):

ok so it is 70.54 then

OpenStudy (perl):

isn't statistics fun :)

OpenStudy (perl):

let me draw what we are doing

OpenStudy (hockeychick23):

hahaha, that was the answer i got in the beginning!

OpenStudy (perl):

yes

OpenStudy (perl):